Find the slope of the line passing through the points (1,1) and (3,3).
Found 2 solutions by checkley71, fractalier:
Answer by checkley71(8403)

SLOPE IS (Y2-Y1)/(X2-X1) OR (3-1)/3-1) OR 2/2 OR 1 IS THE SLOPE.

Answer by fractalier(2101)

Use the formula for the slope, m, of a line:
m = (y2 - y1) / (x2 - x1)
now plug in and go...
m = (3 - 1) / (3 - 1) =
m = 2 / 2
m = 1